Service Status
All Systems Operational
Customer Portals, Tools & Contact Operational
Customer Portal   Operational
Elite's Website   Operational
Ethernet Graphing & Monitoring Portal   ? Operational
Telephone Switchboard   Operational
Ethernet Services Operational
Carrier Interconnects   ? Operational
Broadband Services Operational
WBC ADSL/FTTC Services   Operational
Telehouse North, London Operational
Core Routing   Operational
Ethernet Switching   Operational
Colocation Services   Operational
DNS Services   Operational
Hosting Services   Operational
Telcity HEX, London Operational
Core Routing   Operational
Ethernet Switching   Operational
DNS Services   Operational
Telehouse Metro, London Operational
Core Routing   Operational
Ethernet Switching   Operational
Colocation Services   Operational
Hosting Services   Operational
Backup Services   Operational
Volta GSS, London Operational
Core Routing   Operational
Ethernet Switching   Operational
Colocation Services   Operational
Hosting Services   Operational
Internet House, Portsmouth Operational
Ethernet Switching   Operational
Colocation Services   Operational
DNS Services   Operational
Hosting Services   Operational
Email Services   Operational
Backup Services   Operational
SIP Services Operational
SIP Services   Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Dec 11, 2017

No incidents reported today.

Dec 10, 2017

No incidents reported.

Dec 9, 2017

No incidents reported.

Dec 8, 2017

No incidents reported.

Dec 7, 2017

No incidents reported.

Dec 6, 2017

No incidents reported.

Dec 5, 2017

No incidents reported.

Dec 4, 2017

No incidents reported.

Dec 3, 2017

No incidents reported.

Dec 2, 2017

No incidents reported.

Dec 1, 2017

No incidents reported.

Nov 30, 2017

No incidents reported.

Nov 29, 2017

No incidents reported.

Nov 28, 2017

No incidents reported.

Nov 27, 2017
Resolved - We have observed no further outages to the affected circuits, we will continue to monitor but are happy to mark this resolved.
Nov 27, 13:28 GMT
Monitoring - Affected ethernet circuits are restoring as of 08:59, the tail provider is still investigating the cause of this outage.
Nov 27, 09:37 GMT
Investigating - We are currently investigating an issue affecting multiple leased lines delivered over an NNI, updates are to follow.
Nov 27, 09:03 GMT